# The Unbounded Denominators Conjecture

### v2

@@ 4.2. Modular Forms
    $\ker(f_{1})=G\cap\Gamma_{0}(p)$.
@@
    $\ker(f_{2})=A^{-1}GA\cap\Gamma_{0}(p)$.
Let $K=\langle\ker(f_{1}),\ker(f_{2})\rangle$ be the group generated by $\ker(f_{1})$ and $\ker(f_{2})$. Since $(f_{1},f_{2})$ is surjective, the image of $(f_{1},f_{2})$ contains the elements $(x,0)$ and $(0,y)$ for any $x\in S$. But the pre-images of these elements clearly lie in $\ker(f_{2})$ and $\ker(f_{1})$ respectively, and thus lie in $K$. But then the pre-image of any element lies in $K$, and we deduce that $K=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, or equivalently that
(4.3.10) $\langle G\cap\Gamma_{0}(p),A^{-1}GA\cap\Gamma_{0}(p)\rangle=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p).$
We now complete the proof of Proposition 4.3.1 and hence the proof of Theorem 4.2.6 (as explained at the beginning of § 4.3).
Proof of Proposition 4.3.1.
Consider the function $f(\tau/p)$, which is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. But now conjugating by $A^{-1}$ (to bring $f(\tau/p)$ back to $f(\tau))$ we deduce from Lemma 4.3.2 that $f(\tau)$ is invariant under the group
$A^{-1}(G\cap\Gamma^{0}(p))A=A^{-1}GA\cap\Gamma_{0}(p).$
But by assumption $f(\tau)$ is also invariant under $G$, and hence tautologically under $G\cap\Gamma_{0}(p)$. By equation (4.3.10), $f(\tau)$ is thus also invariant under $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, which is to say invariant under a congruence subgroup. ∎
@@ 5. The uniformization of $\mathbf{C}\smallsetminus\mu_{N}$
We denote by $\mathbf{H}$ the upper half plane and by $\mathbf{P}^{1}=\mathbf{C}\cup\{\infty\}$ the complex projective line or Riemann sphere. There is a conformal isomorphism from the disc $D(0,1)$ to $\mathbf{H}$ by the formula
$x\mapsto i\cdot\frac{1+x}{1-x}.$
@@
Unless we expressly state otherwise, we reserve $z,\tau,x$ to denote the coordinates on $\mathbf{C}\smallsetminus\mu_{N}$ (or $\mathbf{C}\smallsetminus\{1\}$), $\mathbf{H}$, and $D(0,1)$ respectively.
@@ 5.1. Schwarzians and the conformal radius
$F_{N}(x)=\widetilde{F}_{N}\left(i\cdot\frac{1+x}{1-x}\right).$
@@
The action of $\mathrm{SO}_{2}(\mathbf{R})$ on $D(0,1)$ under the pullback map is just given by rotation, and hence $\widetilde{r}_{N}$ acts on $D(0,1)$ by a rotation of order $N$. We deduce that $F_{N}(\zeta^{m}x)=\zeta F_{N}(x)$ for some $(m,N)=1$. By taking the derivatives with respect to $q$ of both sides at $q=0$, we have $\zeta^{m}F_{N}^{\prime}(0)=\zeta F_{N}^{\prime}(0)$. Since $F_{N}$ is a covering map, we must also have $F^{\prime}_{N}(0)\neq 0$, and thus $m=1$ and $F_{N}(\zeta x)=\zeta F_{N}(x)$. We thus also deduce (5.1.3) as this action corresponds to the rotation by degree $2\pi/N$ on $D(0,1)$. ∎
@@
(5.1.11) $\{\tau,1/\widetilde{F}_{N}\}=\frac{1}{2}\sum_{k=0}^{N}\frac{1}{(Z-p_{k})^{2}}+\sum_{k=0}^{N}\frac{m_{k}}{Z-p_{k}},$
where the $m_{k}$ for $k=0,\ldots,N$ denote the so-called accessory parameters at $z=p_{k}$. The accessory parameters are notoriously hard to compute in general, but in our particular example we may find them using the $\mathbf{Z}/N\mathbf{Z}$ symmetry. Expressing the fact that (5.1.11) vanishes to order four at $Z=\infty$, the accessory parameters are subject the following three constraints [Hem88, Theorem 3.1] obtained by equating the $1/Z,1/Z^{2}$ and $1/Z^{3}$ coefficients to zero:
@@
and hence $c=-\frac{1}{2}-\frac{1}{2N}$. This determines all the $m_{k}$, and turns (5.1.11) (still with $Z=1/\widetilde{F}_{N}$) into
$\{\tau,1/\widetilde{F}_{N}\}=\frac{1}{2Z^{2}}+\frac{1}{2}\sum_{k=1}^{N}\frac{1}{(Z-\zeta_{N}^{k})^{2}}-\frac{(1+N)}{2N}\sum_{k=1}^{N}\frac{\zeta_{N}^{-k}}{Z-\zeta_{N}^{k}}=\frac{(1+(N^{2}-1)Z^{N})}{2Z^{2}(Z^{N}-1)^{2}}.$
From the chain rule, we deduce that with $z=\widetilde{F}_{N}=1/Z$ the equality:
$\{\tau,\widetilde{F}_{N}\}=\frac{1}{z^{4}}\frac{(1+(N^{2}-1)(1/z)^{N})}{2(1/z)^{2}((1/z)^{N}-1)^{2}}=\frac{(N^{2}-1)z^{N-2}+z^{2N-2}}{2(z^{N}-1)^{2}},$
@@
Let $G_{N}$ denote the map $D(0,1)\rightarrow\mathbf{C}\smallsetminus\{1\}$ such that $G_{N}(x^{N})=(F_{N}(x))^{N}$, or equivalently $G_{N}(x)=(F_{N}(x^{1/N}))^{N}$.
The fact that $G_{N}$ is well-defined is a formal consequence of the relation $F_{N}(\zeta x)=\zeta F_{N}(x)$ in Lemma 5.1.2.
@@
Let $\varphi_{N}$ denote the local inverse map of $G_{N}$ around $z=0$, normalized so that $\varphi_{N}(0)=0$. The function $\varphi_{N}$ has the form $\delta^{-1}_{N}(\phi_{1}/\phi_{2})^{N}$, where $\phi_{1}$ and $\phi_{2}$ are the solutions to the differential equation:
(5.1.15) $z(z-1)^{2}y^{\prime\prime}+\left(1-\frac{1}{N}\right)(z-1)^{2}y^{\prime}+\left(\frac{1}{4}+\frac{z-1}{4N^{2}}\right)y=0$
@@
Moreover, let $s_{N}(z)$ denote the function
@@
Then $\varphi_{N}(z)=\delta_{N}^{-1}s_{N}(z)^{N}$, $\psi_{N}(z)=|F^{\prime}_{N}(0)|^{-1}s_{N}(z^{N})$, and $\delta_{N}=|F^{\prime}_{N}(0)|^{N}$.
@@
By Definition 5.1.13, $G_{N}(x)=(F_{N}(x^{1/N}))^{N}$; and by the assumptions $\varphi_{N}(0)=0$ and $\psi_{N}(0)=0$, we obtain the formal identity $\varphi_{N}(z)=\psi_{N}(z^{1/N})^{N}$ (here formally take $z=G_{N}(x)$).
Let $\eta_{1},\eta_{2}$ denote the solutions of the differential equation in Lemma 5.1.8 such that $\eta_{1}(0)=0,\eta^{\prime}_{1}(0)=1,\eta_{2}(0)=1,\eta^{\prime}_{2}(0)=0$; then $\eta_{1},\eta_{2}$ are linearly independent and $\eta_{1}(z)/\eta_{2}(z)=z+O(z^{2})$. Since $\psi_{N}(z)=|F_{N}^{\prime}(0)|^{-1}z+O(z^{2})$, then by Lemma 5.1.8, we have $\psi_{N}=|F_{N}^{\prime}(0)|^{-1}\,\eta_{1}/\eta_{2}$. We deduce
$\varphi_{N}(z)=|F_{N}^{\prime}(0)|^{-N}(\eta_{1}(z^{1/N})/\eta_{2}(z^{1/N}))^{N}.$
Let $\phi_{i}(z)=\eta_{i}(z^{1/N})$. Then
$\phi^{\prime}_{i}(z)=N^{-1}z^{\frac{1}{N}-1}\eta^{\prime}_{i}(z^{1/N})$
@@
$\phi^{\prime\prime}_{i}(z)=\frac{1-N}{N^{2}}z^{\frac{1}{N}-2}\eta^{\prime}_{i}+N^{-2}z^{\frac{2}{N}-2}\eta^{\prime\prime}_{i}(z^{1/N}).$
@@
$4(z-1)^{2}\eta^{\prime\prime}_{i}(z^{1/N})+((N^{2}-1)z^{1-\frac{2}{N}}+z^{2-\frac{2}{N}})\eta_{i}(z^{1/N})=0.$
@@
In order to prove Theorem 5.1.4, we need the following formula of the behavior of hypergeometric functions in Lemma 5.1.14 near $z=1$.
@@
Given $a\notin\mathbf{Z}_{\leq 0}$, for $|z|<1$, we have
@@
where $\psi$ denotes the logarithmic derivative of the Gamma function. Note that the above hypergeometric function is multivalued around $z=0$, but all different branches are accounted for by the branches of the logarithm. ∎
@@
Since $F_{N}$ is a covering map of $\mathbf{C}\smallsetminus\mu_{N}$, then local inverse $\psi_{N}$ is naturally defined on $D(0,1)\subset\mathbf{C}\smallsetminus\mu_{N}$. Moreover, since $F_{N}(1)=1$, we have that $\lim_{z\rightarrow 1}\psi_{N}(z)=1$, where $z\in D(0,1)$ approaches $1$. (A priori, we could only conclude that $\lim_{z\rightarrow 1}\psi_{N}(z)$ approaches a cusp of $D(0,1)$ and thus $|\lim_{z\rightarrow 1}\psi_{N}(z)|=1$, which is enough for the rest of the proof; on the other hand, either by the description of the fundamental domain in Lemma 5.2.1 or by the computation of the rest of the proof that $\lim_{z\rightarrow 1}\psi_{N}(z)$ is a positive real number, we indeed have $\lim_{z\rightarrow 1}\psi_{N}(z)=1$ due to our assumption that $\psi_{N}(0)=0$.)
By Lemma 5.1.14, we have $\psi_{N}(z)=|F^{\prime}_{N}(0)|^{-1}s_{N}(z^{N})$. In particular,
$|F^{\prime}_{N}(0)|^{-1}\lim_{z\in D(0,1),z\rightarrow 1}s_{N}(z^{N})=1.$

### v3

@@ 4.5. Invariant vectors
    For $N=3$, we have $\mathrm{PSL}_{2}(\mathbf{F}_{3})=A_{4}$, and $V_{\mathbf{Q}}$ is absolutely irreducible of dimension $3$. The associated Brauer character is also irreducible and so $V_{\mathbf{F}_{\ell}}]$ is also irreducible and has no invariants.
@@
    For $N=5$, we have $\mathrm{PSL}_{2}(\mathbf{Z}/5\mathbf{Z})\simeq A_{5}$, and $V_{\mathbf{Q}}$ decomposes as a sum of irreducibles of dimensions $3$, 33, and $5$. The corresponding Brauer characters are all still irreducible, so $V_{\mathbf{F}_{\ell}}]$ does not contain the trivial representation.
  3. (3)
    For $N=4$, we have $\mathrm{PSL}_{2}(\mathbf{Z}/4\mathbf{Z})=S_{4}$, and $V_{\mathbf{Q}}$ is a sum of absolutely irreducible representations of dimensions $2$ and $3$. The group $S_{4}$ has two Brauer characters of dimension $1$ and $2$ respectively. The $2$ dimensional representation remains irreducible and the semi-simplification of both the $3$ dimensional representations has constituents of dimensions $1$ and $2$. Hence the dimension of the invariant space of $V^{\mathrm{ss}}_{\mathbf{F}_{\ell}}$ is $1$-dimensional. But $H^{1}(\Gamma(4)/\Gamma(8),\mathbf{F}_{2})$ is a direct sum of the $1$ and $2$ dimensional representations, so this one-dimensional constituent occurs as a sub-representation.
  4. (4)
    For $N=8$ and $N=16$, we resort to a less elegant calculation; the groups $\Gamma(N)$ are free (of ranks $33$ and $257$ respectively). The $\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$-invariant part of cohomology over $\mathbf{F}_{2}$ can be determined as (the dual of) the quotient of this group by the relations $x^{2}=(xy)^{2}=e$ for each generator $x\in\Gamma(N)$ and the relations $gxg^{-1}=x$ for the generators $g$ of $\mathrm{SL}_{2}(\mathbf{Z})$. In both cases, magma determines that the corresponding quotients have order $2$.
Proof of Theorem 4.5.2.
We now complete the proof of Theorem 4.5.2.
Let $v\in\widetilde{H}^{1}(\mathbf{F}_{\ell})^{\mathrm{SL}_{2}(\widehat{\mathbf{Z}})}$. From the Künneth formula and Lemma 4.5.6, we deduce that there is an isomorphism $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})=0$ for any prime $\ell>5$. Thus, for $\ell>5$, any class $v$ lies in the image of $H^{1}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{\ell})\simeq\mathbf{Z}/12\mathbf{Z}\otimes\mathbf{F}_{\ell}$, but one easily checks that the map
$H^{1}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})\rightarrow H^{1}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{\ell})$
is an isomorphism, and thus $v=0$.
Suppose that $\ell=3$ or $\ell=5$. Since the map $H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{\ell})\rightarrow H^{2}(\widehat{\Gamma}(\ell),\mathbf{F}_{\ell})$ is zero by Lemma 4.5.9, the image of $v$ lands in the $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-invariants of $H^{1}(\Gamma(\ell),\mathbf{F}_{\ell})$. Thus it contributes to the $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-invariants of the semi-simplification of $H^{1}(\Gamma(\ell),\mathbf{F}_{\ell})$ as a $\mathrm{SL}_{2}(\mathbf{Z}/\ell\mathbf{Z})$-module. But this space has dimension zero by Lemma 4.5.10.
Finally, let $\ell=2$. By Lemma 4.5.9, the map
$H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{2})\rightarrow H^{2}(\widehat{\Gamma}(8),\mathbf{F}_{2})$
is zero, and thus the element $v$ lies in the image of $H^{1}(\Gamma(8),\mathbf{F}_{2})$. Furthermore, it has the property that the $\mathrm{SL}_{2}(\mathbf{Z}/8\mathbf{Z})$-module generated by $v$ is trivial after passing to the quotient by the congruence homology
$H^{1}(\widehat{\Gamma}(8),\mathbf{F}_{2})\simeq H^{1}(\widehat{\Gamma}(8)/\widehat{\Gamma}(16),\mathbf{F}_{2}).$
But that means that the image of $v$ in $H^{1}(\widehat{\Gamma}(16),\mathbf{F}_{2})$ is genuinely invariant under $\mathrm{SL}_{2}(\mathbf{Z}/16\mathbf{Z})$. By Lemma 4.5.10, the space of such invariants is one dimensional. But this one dimensional space lands in the image of $H^{1}(\widehat{\Gamma}(16),\mathbf{F}_{2})$, and thus $v$ is trivial in $\widetilde{H}^{1}(\mathbf{F}_{2})$. ∎
We note in passing that this result implies the following strengthening of Lemma 4.5.6:
Lemma 4.5.11.
For $n=1$ and $n=2$ we have:
$H^{n}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})=\begin{cases}\mathbf{F}_{\ell},&p=\ell\in\{2,3\}\\ 0,&\text{otherwise}.\end{cases}$
We also have $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})=0$ for all $p$.
Proof.
From Theorem 4.5.6, it suffices to consider the case of $n=2$ and $\ell=p\in\{2,3,5\}$. There is an exact sequence:
$0\rightarrow H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})/\ell\rightarrow H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{\ell})\rightarrow H_{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})[\ell]\rightarrow 0.$
Since $H_{1}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})\simeq\mathbf{Z}/12\mathbf{Z}\otimes\mathbf{Z}_{p}$, this implies the desired lower bounds for in Lemma 4.5.11, with equality in all cases if and only if $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})/\ell=0$ for all primes $\ell$. Since the group $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})$ is a finitely generated abelian group, to deduce that $H_{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{Z})=0$ it suffices to prove the upper bounds.
For the upper bounds, we deduce from Theorem 4.5.2 that there is an injection
$H^{2}(\mathrm{SL}_{2}(\mathbf{Z}_{p}),\mathbf{F}_{p})\rightarrow H^{2}(\mathrm{SL}_{2}(\widehat{\mathbf{Z}}),\mathbf{F}_{p})\rightarrow H^{2}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{p}).$
But for $n>0$ we have
$H^{n}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{Z})=\begin{cases}\mathbf{Z}/12\mathbf{Z},&n\equiv 0\bmod 2,\\ 0,&n\equiv 1\bmod 2\end{cases},$
from which it follows that $H^{2}(\mathrm{SL}_{2}(\mathbf{Z}),\mathbf{F}_{p})=\mathbf{F}_{p}$ if $p\in\{2,3\}$ and is zero otherwise. This completes the proof. ∎
4.6. An enhancement of Ihara’s Lemma
We shall prove an enhanced version of Ihara’s Lemma. We begin by recalling Ihara’s Lemma. Let $q$ be prime, let $N\geq 3$, and let $(N,p)=1$. There is a homomorphism
(4.6.1) $H^{1}(\Gamma(N),\mathbf{F}_{q})^{\oplus 2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q}),$
given by the difference of the following two maps:
  1. (1)
    The map sending $\psi:\Gamma(N)\rightarrow\mathbf{F}_{q}$ to its restriction to $\Gamma(N)\cap\Gamma_{0}(p)$.
  2. (2)
    The twisted restriction map coming from viewing $\psi\in H^{1}(\Gamma(N),\mathbf{F}_{q})$ as a map from $\Gamma(N)\rightarrow\mathbf{F}_{q}$ and then considering the map
    $A\psi:\Gamma(N)\cap\Gamma_{0}(p),\quad g\mapsto\psi(AgA^{-1}).$
By abuse of notation we denote the restriction of $\psi$ by $\psi$, so the map sends $(\psi,\phi)$ to $\psi-A\phi$.
Lemma 4.6.2 (Ihara’s Lemma).
The kernel of the map (4.6.1) lies inside $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{q})^{2}$.
Proof.
This version of Ihara’s Lemma was essentially proved by Ribet in [Rib84]. The proof is just an abelian version of Lemma 4.4.1. We recall some of the details. Let $\Phi\subset\Gamma(N)$ be the maximal normal subgroup whose quotient is an elementary $q$-abelian group $T$. Canonically, we have $\Gamma(N)/\Phi\simeq T$ and $H^{1}(\Gamma,\mathbf{F}_{q})\simeq\mathrm{Hom}(T,\mathbf{F}_{q})$. The kernel of the map
$\mathrm{Hom}(T,\mathbf{F}_{q})\times\mathrm{Hom}(T,\mathbf{F}_{q})\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q})$
is governed by the cokernel of the dual map
$\Gamma(N)\cap\Gamma_{0}(p)\rightarrow T\times T.$
Exactly as in the proof of Lemma 4.4.1, we deduce from Goursat’s Lemma that the cokernel $\Delta$ arises from two maps from $\Gamma(N)$ to $\Delta$ which agree along $\Gamma(N)\cap\Gamma_{0}(p)$, and thus on their amalgam $\mathrm{SL}_{2}(\mathbf{Z}[1/p])(N)$. Since $\mathrm{SL}_{2}(\mathbf{Z}[1/p])(N)$ as the congruence subgroup property, it thus arises from a congruence quotient of this group at primes away from $p$. But that precisely means that the classes in $\mathrm{Hom}(T,\mathbf{F}_{q})=H^{1}(\Gamma(N),\mathbf{F}_{q})$ become trivial after passing to a congruence subgroup $\Gamma^{\prime}\subset\Gamma(N)$, hence the claim. ∎
Using Corollary 4.5.3, we prove a slight enhancement of this claim.
Lemma 4.6.3 (Ihara’s Lemma, enhanced).
The kernel of the composite of the Ihara map (4.6.1) with the map
(4.6.4) $H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q})\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q})$
also lies inside $H^{1,\mathrm{cong}}(\Gamma,\mathbf{F}_{q})^{2}$.
Proof.
The map (4.6.1) is $\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$-equivariant. But the kernel of the map (4.6.4) is also easily seen to be $\mathrm{SL}_{2}(\mathbf{Z}/N\mathbf{Z})$-invariant. Hence, if $(\psi,\phi)$ lies in the kernel of the composite of (4.6.1) and (4.6.4), then $(\psi^{g}-\psi,\phi^{g}-\phi)$ lies in the kernel of (4.6.1), and thus lies in $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})^{2}$ by Lemma 4.6.2. But then $\psi$ and $\phi$ are themselves congruence classes by Corollary 4.5.3. ∎
@@ 5. The uniformization of $\mathbf{C}\smallsetminus\mu_{N}$
We denote by $\mathbf{H}$ the upper half plane and by $\mathbf{P}^{1}=\mathbf{C}\cup\{\infty\}$ the complex projective line or Riemann sphere. There is a conformal isomorphism from the disc $D(0,1)$ to $\mathbf{H}$ by the Cayley transform
$z\mapsto i\cdot\frac{1+z}{1-z}.$
@@
Unless we expressly state otherwise, we reserve $z,\tau,x$ to denote respectively the coordinates on $D(0,1)$, $\mathbf{H}$, and $\mathbf{C}\smallsetminus\mu_{N}$.
@@ 5.1. Schwarzians and the conformal radius
$F_{N}(z)=\widetilde{F}_{N}\left(i\cdot\frac{1+z}{1-z}\right).$
@@
The action of $\mathrm{SO}_{2}(\mathbf{R})$ on $D(0,1)$ under the pullback map is just given by rotation, and hence $\widetilde{r}_{N}$ acts on $D(0,1)$ by a rotation of order $N$. We deduce that $F_{N}(\zeta^{m}z)=\zeta F_{N}(z)$ for some $(m,N)=1$. By taking the derivatives with respect to $q$ of both sides at $q=0$, we have $\zeta^{m}F_{N}^{\prime}(0)=\zeta F_{N}^{\prime}(0)$. Since $F_{N}$ is a covering map, we must also have $F^{\prime}_{N}(0)\neq 0$, and thus $m=1$ and $F_{N}(\zeta z)=\zeta F_{N}(z)$. We thus also deduce (5.1.3) as this action corresponds to the rotation by degree $2\pi/N$ on $D(0,1)$. ∎
@@
(5.1.11) $\{\tau,1/\widetilde{F}_{N}\}=\frac{1}{2}\sum_{k=0}^{N}\frac{1}{(X-p_{k})^{2}}+\sum_{k=0}^{N}\frac{m_{k}}{X-p_{k}},$
where the $m_{k}$ for $k=0,\ldots,N$ denote the so-called accessory parameters at $z=p_{k}$. The accessory parameters are notoriously hard to compute in general, but in our particular example we may find them using the $\mathbf{Z}/N\mathbf{Z}$ symmetry. Expressing the fact that (5.1.11) vanishes to order four at $X=\infty$, the accessory parameters are subject to the following three constraints [Hem88, Theorem 3.1] obtained by equating the $1/X,1/X^{2}$ and $1/X^{3}$ coefficients to zero:
@@
and hence $c=-\frac{1}{2}-\frac{1}{2N}$. This determines all the $m_{k}$, and turns (5.1.11) (still with $X=1/\widetilde{F}_{N}$) into
$\{\tau,1/\widetilde{F}_{N}\}=\frac{1}{2X^{2}}+\frac{1}{2}\sum_{k=1}^{N}\frac{1}{(X-\zeta_{N}^{k})^{2}}-\frac{(1+N)}{2N}\sum_{k=1}^{N}\frac{\zeta_{N}^{-k}}{X-\zeta_{N}^{k}}=\frac{(1+(N^{2}-1)X^{N})}{2X^{2}(X^{N}-1)^{2}}.$
From the chain rule, we deduce that with $x=\widetilde{F}_{N}=1/X$ the equality:
$\{\tau,\widetilde{F}_{N}\}=\frac{1}{x^{4}}\frac{(1+(N^{2}-1)(1/x)^{N})}{2(1/x)^{2}((1/x)^{N}-1)^{2}}=\frac{(N^{2}-1)x^{N-2}+x^{2N-2}}{2(x^{N}-1)^{2}},$
@@
Let $G_{N}$ denote the map $D(0,1)\rightarrow\mathbf{C}\smallsetminus\{1\}$ such that $G_{N}(z^{N})=(F_{N}(z))^{N}$, or equivalently $G_{N}(z)=(F_{N}(z^{1/N}))^{N}$.
The fact that $G_{N}$ is well-defined is a formal consequence of the relation $F_{N}(\zeta z)=\zeta F_{N}(z)$ in Lemma 5.1.2.
@@
Let $\varphi_{N}$ denote the local inverse map of $G_{N}$ around $x=0$, normalized so that $\varphi_{N}(0)=0$. The function $\varphi_{N}$ has the form $\delta^{-1}_{N}(\phi_{1}/\phi_{2})^{N}$, where $\phi_{1}$ and $\phi_{2}$ are the solutions to the differential equation:
(5.1.15) $x(x-1)^{2}y^{\prime\prime}+\left(1-\frac{1}{N}\right)(x-1)^{2}y^{\prime}+\left(\frac{1}{4}+\frac{x-1}{4N^{2}}\right)y=0$
@@
Further, let $s_{N}(x)$ denote the function
@@
Then $\varphi_{N}(x)=\delta_{N}^{-1}s_{N}(x)^{N}$, $\psi_{N}(x)=|F^{\prime}_{N}(0)|^{-1}s_{N}(x^{N})$, and $\delta_{N}=|F^{\prime}_{N}(0)|^{N}$.
@@
By Definition 5.1.13, $G_{N}(z)=(F_{N}(z^{1/N}))^{N}$; and by the assumptions $\varphi_{N}(0)=0$ and $\psi_{N}(0)=0$, we obtain the formal identity $\varphi_{N}(x)=\psi_{N}(x^{1/N})^{N}$ (formally: $x=G_{N}(z)$).
Let $\eta_{1},\eta_{2}$ denote the solutions of the differential equation in Lemma 5.1.8 such that $\eta_{1}(0)=0,\eta^{\prime}_{1}(0)=1,\eta_{2}(0)=1,\eta^{\prime}_{2}(0)=0$; then $\eta_{1},\eta_{2}$ are linearly independent and $\eta_{1}(x)/\eta_{2}(x)=x+O(x^{2})$. Since $\psi_{N}(x)=|F_{N}^{\prime}(0)|^{-1}x+O(x^{2})$, then by Lemma 5.1.8, we have $\psi_{N}=|F_{N}^{\prime}(0)|^{-1}\,\eta_{1}/\eta_{2}$. We deduce
$\varphi_{N}(x)=|F_{N}^{\prime}(0)|^{-N}(\eta_{1}(x^{1/N})/\eta_{2}(x^{1/N}))^{N}.$
Let $\phi_{i}(x)=\eta_{i}(x^{1/N})$. Then
$\phi^{\prime}_{i}(x)=N^{-1}x^{\frac{1}{N}-1}\eta^{\prime}_{i}(x^{1/N})$
@@
$\phi^{\prime\prime}_{i}(x)=\frac{1-N}{N^{2}}x^{\frac{1}{N}-2}\eta^{\prime}_{i}+N^{-2}x^{\frac{2}{N}-2}\eta^{\prime\prime}_{i}(x^{1/N}).$
@@
$4(x-1)^{2}\eta^{\prime\prime}_{i}(x^{1/N})+((N^{2}-1)x^{1-\frac{2}{N}}+x^{2-\frac{2}{N}})\eta_{i}(x^{1/N})=0.$
@@
In order to prove Theorem 5.1.4, we need the following formula of the behavior of hypergeometric functions in Lemma 5.1.14 near $x=1$.
@@
Given $a\notin\mathbf{Z}_{\leq 0}$, for $|x|<1$, we have
@@
where $\psi$ denotes the digamma function: the logarithmic derivative of the Gamma function. Note that the above hypergeometric function is multivalued around $x=0$, but all different branches are accounted for by the branches of the logarithm. ∎
@@
Since $F_{N}$ is a covering map of $\mathbf{C}\smallsetminus\mu_{N}$, then local inverse $\psi_{N}$ is naturally defined on $D(0,1)\subset\mathbf{C}\smallsetminus\mu_{N}$. Moreover, since $F_{N}(1)=1$, we have that $\lim_{x\rightarrow 1}\psi_{N}(x)=1$, where $x\in D(0,1)$ approaches $1$. (A priori, we only conclude that $\lim_{x\rightarrow 1}\psi_{N}(x)$ approaches a cusp of $D(0,1)$, and thus that $|\lim_{x\rightarrow 1}\psi_{N}(x)|=1$; this suffices for the rest of the proof. The more precise statement $\lim_{x\rightarrow 1}\psi_{N}(x)=1$ follows from our assumption $\psi_{N}(0)=0$, either by the description of the fundamental domain further down in Lemma 5.2.1, or more directly by the computation of the rest of the proof that follows, which shows that that $\lim_{x\rightarrow 1}\psi_{N}(x)$ is a positive real number.)
By Lemma 5.1.14, we have $\psi_{N}(x)=|F^{\prime}_{N}(0)|^{-1}s_{N}(x^{N})$. In particular,
$|F^{\prime}_{N}(0)|^{-1}\lim_{x\in D(0,1),x\rightarrow 1}s_{N}(x^{N})=1.$